# UniGraph2: Learning a Unified Embedding Space to Bind Multimodal Graphs

### v1

@@ 1. Introduction
Real-world web applications increasingly rely on multimodal data, where information is derived from a variety of sources such as text, images, and audio (Ngiam et al., 2011; Baltrušaitis et al., 2018). Recent foundation models have focused on learning a unified embedding space across different modalities that allows for the seamless integration of multimodal data, thereby enabling effective cross-modal interactions and supporting downstream applications (Radford et al., 2021; Girdhar et al., 2023).
Models such as CLIP (Radford et al., 2021) have demonstrated the power of learning from multimodal data by mapping text and images into a shared embedding space. However, CLIP and similar models are fundamentally limited by their reliance on a 1-to-1 mapping between paired modalities, such as text-to-image alignment, ignoring more complex structures where nodes can be connected through many-to-many relationships and involve multiple modalities. These models fail to account for the graph structure present in numerous web domains, from social networks to e-commerce networks (Yoon et al., 2023; Zhu et al., 2024; Ektefaie et al., 2023), where entities and their interactions are crucial to understanding the underlying relationships. For example, in e-commerce platforms, recommendation systems rely on complex networks of products, users, and their interactions (Schafer et al., 2001). Each node represents a a user or a product, and edges represent interactions like purchases, views, or reviews. Additionally, both users and products are associated with rich multimodal data: product descriptions (text), images (visual), and user reviews (text), and demonstration videos (audio and visual). Integrating these diverse data types within the graph structure is essential for accurate recommendations and personalized user experiences (Gao et al., 2022). To address these challenges, Multimodal Graphs (MMGs) have been introduced as a framework that combines graph structures with multimodal data (Zhu et al., 2024; Ektefaie et al., 2023). On MMGs, nodes are enriched with information from multiple modalities, allowing for a more comprehensive representation of entities and their relationships. However, existing MMGs learning methods can only train models individually for a specific graph and task (Yoon et al., 2023; Chen et al., 2022; Zeng et al., 2023), and cannot achieve cross-graph and cross-task transfer like foundation models do without retraining or fine-tuning.
@@ 2.2. Multimodal Graph Learning
Most existing multimodal graph learning models primarily focus on knowledge graphs (Chen et al., 2022; Zeng et al., 2023) and natural sciences, such as molecular graphs (Jin et al., 2019) or brain graphs (Wang et al., 2023a). However, these models are specifically designed for particular tasks on individual graphs using domain knowledge and do not aim to learn a unified and general representation. They also cannot be transferred across different graphs, modalities, or tasks. Unlike these works, a recent work, MMGL (Yoon et al., 2023) explores the use of foundation models from different modalities on MMGs, but it focuses solely on generative tasks.
@@ 2.3. Graph Foundation Models
Learning graph foundation models that can be transferred across different graphs (He et al., 2024d, c; Qiu et al., 2020; Hou et al., 2024) and tasks (Hou et al., 2023; Liu et al., 2024a; He et al., 2024c) has recently received significant attention. Some works explore designing domain-specific graph foundation models, such as those for knowledge graphs (Galkin et al., 2024) and molecular graphs (Xia et al., 2023). Most existing research efforts are dedicated to using LLMs with strong generalization capabilities to solve graph learning tasks (Wang et al., 2024a; He et al., 2024e; Sui et al., 2024; Liu et al., 2024a). However, how to effectively serialize graph data so that LLMs can understand the graph structure and graph learning tasks remains a barrier to further performance improvements (Zhang et al., 2024). Additionally, these models typically use the generative capabilities of LLMs to directly generate predicted labels, thus addressing representation learning tasks on graphs. Due to the high computational cost, it is challenging to scale them to web-scale large graphs (Wang et al., 2024a; He et al., 2024e).

### v2

@@ 1. Introduction
Real-world web applications increasingly rely on multimodal data, where information is derived from a variety of sources such as text, images, and audio (Ngiam et al., 2011; Baltrušaitis et al., 2018; Sui et al., 2025; He et al., 2025; Chen et al., 2025; Liu et al., 2025). Recent foundation models have focused on learning a unified embedding space across different modalities that allows for the seamless integration of multimodal data, thereby enabling effective cross-modal interactions and supporting downstream applications (Radford et al., 2021; Girdhar et al., 2023; Gao et al., 2025).
Models such as CLIP (Radford et al., 2021) have demonstrated the power of learning from multimodal data by mapping text and images into a shared embedding space. However, CLIP and similar models are fundamentally limited by their reliance on a 1-to-1 mapping between paired modalities, such as text-to-image alignment, ignoring more complex structures where nodes can be connected through many-to-many relationships and involve multiple modalities. These models fail to account for the graph structure present in numerous web domains, from social networks to e-commerce networks (Yoon et al., 2023; Zhu et al., 2024; Ektefaie et al., 2023; Zhang et al., 2021; He and Ma, 2022), where entities and their interactions are crucial to understanding the underlying relationships. For example, in e-commerce platforms, recommendation systems rely on complex networks of products, users, and their interactions (Schafer et al., 2001). Each node represents a a user or a product, and edges represent interactions like purchases, views, or reviews. Additionally, both users and products are associated with rich multimodal data: product descriptions (text), images (visual), and user reviews (text), and demonstration videos (audio and visual). Integrating these diverse data types within the graph structure is essential for accurate recommendations and personalized user experiences (Gao et al., 2022). To address these challenges, Multimodal Graphs (MMGs) have been introduced as a framework that combines graph structures with multimodal data (Zhu et al., 2024; Ektefaie et al., 2023). On MMGs, nodes are enriched with information from multiple modalities, allowing for a more comprehensive representation of entities and their relationships. However, existing MMGs learning methods can only train models individually for a specific graph and task (Yoon et al., 2023; Chen et al., 2022; Zeng et al., 2023), and cannot achieve cross-graph and cross-task transfer like foundation models do without retraining or fine-tuning.
@@ 2.2. Multimodal Graph Learning
Most existing multimodal graph learning models primarily focus on knowledge graphs (Chen et al., 2022; Zeng et al., 2023; Sui et al., 2024a) and natural sciences, such as molecular graphs (Jin et al., 2019) or brain graphs (Wang et al., 2023a). However, these models are specifically designed for particular tasks on individual graphs using domain knowledge and do not aim to learn a unified and general representation. They also cannot be transferred across different graphs, modalities, or tasks. Unlike these works, a recent work, MMGL (Yoon et al., 2023) explores the use of foundation models from different modalities on MMGs, but it focuses solely on generative tasks.
@@ 2.3. Graph Foundation Models
Learning graph foundation models that can be transferred across different graphs (He et al., 2024d, c; Qiu et al., 2020; Hou et al., 2024) and tasks (Hou et al., 2023; Liu et al., 2024a; He et al., 2024c, 2022) has recently received significant attention. Some works explore designing domain-specific graph foundation models, such as those for knowledge graphs (Galkin et al., 2024; Sui et al., 2024b) and molecular graphs (Xia et al., 2023). Most existing research efforts are dedicated to using LLMs with strong generalization capabilities to solve graph learning tasks (Wang et al., 2024a; He et al., 2024e; Sui et al., 2024b; Liu et al., 2024a). However, how to effectively serialize graph data so that LLMs can understand the graph structure and graph learning tasks remains a barrier to further performance improvements (Zhang et al., 2024). Additionally, these models typically use the generative capabilities of LLMs to directly generate predicted labels, thus addressing representation learning tasks on graphs. Due to the high computational cost, it is challenging to scale them to web-scale large graphs (Wang et al., 2024a; He et al., 2024e).